# Demonstration of Efficient Radon Removal by Silver-Zeolite in a Dark Matter Detector

### v1

@@ II.1 Quality cut
We apply a quality cut that includes a loose pulse shape discrimination cut to remove non-physical events with characteristically short risetimes [72], as well as requiring pulse amplitudes above 1200 ADU (arbitrary digital units). The latter selection removes most of the background events from atmospheric muons, as well as additional electronic noise events. Some degraded alpha events that do not deposit their full kinetic energy also fall below this cut. However, the fraction of affected alpha events remains mostly constant over time, tending to increase slightly as the detector gain decreases. A separate study quantifying the impact of this effect on signal efficiency due to this cut shows a drop of only $\sim$ 1% over time (see Fig. 3.59 in Ref. [90]).
@@
Figure S1: Amplitude distribution of events for approximately 6 hours of data in phase III recorded by the SPC (black markers). The distribution is modeled with an adaptive-bandwidth Gaussian kernel density estimation [101, 102, 103, 104] with bootstrapped [105] statistical uncertainty (red curve and $1\sigma$ shaded band). The inset shows the amplitude peak positions corresponding to the maximum energy deposition of three alpha decays versus the corresponding alpha energies (red markers with error bars), with a linear fit (blue curve and shaded band).
@@
The remaining constant background in phase I for all campaigns (see Fig. 2 in the main text) comes predominantly from the alpha decay of 210Po (5.3 MeV [87]), originating from radioisotope depositions on the inner surface of the detector material. The SPC, made of stainless steel, does not have particular precautions to prevent 210Po contamination, but can nonetheless serve as a useful low-rate calibration source for this analysis. In phase IV, following the opening of the trap, the rate in campaigns 1 and 2 decreases significantly over several hours, eventually reaching the 210Po alpha background level observed in phase I. For campaign 3, using activated charcoal, the rate remains up to two orders of magnitude higher compared to the rate in phase I and exhibits exponential decay, indicating the continued presence of radon.
@@
In phases II and III, with the injection of radon, the rate increases. By applying the quality cut, we select primarily the full energy alpha decays of  [^2] Rn (5.5 MeV [40]) and its daughter isotopes, 218Po (6.0 MeV [88]) and 214Po (7.7 MeV [89]), which have clear amplitude peaks (above 12,000 ADU) in the recorded data, as shown in Fig. S1 for phase III. The nominal radon trap analysis described in the main text is performed on the data with only this quality cut applied, incorporating events of all three alpha decay isotopes; as these decays are in secular equilibrium, the proportion of events due to radon itself is approximately constant throughout each measurement campaign. Additionally, high energy beta decay events from the  [^2] Rn decay chain (214Bi in particular [88]) are present in this data selection, but are also in equilibrium with the rest of the decay chain.
@@
To verify the linear energy response of the SPC in this regime, amplitude spectra—such as the one shown in Fig. S1—are characterized with an adaptive-bandwidth Gaussian kernel density estimation [101, 102, 103, 104], with statistical uncertainties being estimated using a bootstrapping procedure [105]. The resulting peak amplitudes are well-fit by a linear function of the corresponding alpha decay energies, as shown in the inset of Fig. S1. A peak corresponding to 210Po might be present in this amplitude spectrum, but its rate is more than two orders of magnitude lower than the  [^2] Rn peak with which it overlaps.
@@
In our analysis, we develop a method to select only 214Po events. 214Po decays achieve secular equilibrium with  [^2] Rn due to its short half-life (164 ${\upmu}$s [89]) compared to  [^2] Rn and its subsequent daughter isotopes. By selecting only 214Po events, the signal-to-background ratio is significantly improved, as there is no 210Po background in this amplitude regime.
@@
Due to the significant amount of radon remaining in phase IV of campaign 3 (see Fig. 2 in the main text), the 214Po amplitude cut is determined using a similar method as in phases II and III. To obtain a time-dependent 214Po amplitude cut for phase IV, the amplitude minima between the 218Po and 214Po peaks are found in approximately 6-hour data segments. The bootstrap resampling [105] is then applied to estimate the statistical uncertainty for each point. These points are then fit with a linear function to define the 214Po amplitude cut for phase IV.
@@ II.3 Expected decay rates
To achieve this, a Monte Carlo (MC) simulation of radon decay is produced for a given radon rate in addition to a given background rate, which includes the expected pileup and dead-time effects to reflect the actual conditions of our experiment. Any simulated decays occurring within 1 ms of each other (half the duration of a single event window) are treated as coincident events (resulting in a single recorded event). Simulated decays occurring within 1 to 2 ms of a preceding pulse are considered as lost due to DAQ deadtime. This approach accounts for the DAQ software centering the maximum amplitude of the triggered event trace within its acquisition time window, ensuring that no two pulses can be more than half an event window apart and still be contained in the event window. High-statistics MCs are produced for true radon rates ranging from approximately 30 Hz to 200 Hz, along with the measured 210Po background rate for each campaign described in the main text. Each MC dataset spans a duration longer than the real data from phase III, covering both phase III and phase IV. The resulting log-scale rate trends—separated into single decay event vs. pileup events—are characterized using univariate splines [106].
The results of the decay rate MCs are incorporated into a joint model for the observed event rate over time, with and without the application of the 214Po cut. The free parameters of the model define: (i) the true radon decay rate from the discrete sets of values for which MCs are produced; (ii) the overall alpha event signal efficiency with the quality cut applied ($f_{\alpha}$); (iii) signal efficiency of observed 214Po events with the 214Po cut applied ($f_{\text{Po-214}}$); (iv) the fraction of pileup events in the data without and with the 214Po cut ($f_{\text{pileup}}$ and $f_{\text{pileup, Po-214}}$, respectively). Adding the background rates for each data selection ($\eta$ and $\eta_{\text{Po-214}}$), the expected event rates over time $\nu$, with either the quality or 214Po cut, are given as
$\nu$ $(t)=f_{\alpha}\left[10^{y_{\text{Rn-222}}(t)}+10^{y_{\text{Po-218}}(t)}+10^{y_{\text{Po-214}}(t)}\right]{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\frac{|}{|}}$
$+f_{\text{pileup}}\times 10^{y_{\text{pileup}}(t)}+\eta,{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\frac{|}{|}}$
$\nu$ (${}_{\text{Po-214}}(t)=f_{\text{Po-214}}\times 10^{y_{\text{Po-214}}(t)}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\frac{|}{|}}$
@@
$\log$ $\mathcal{L}\ =\sum_{i}\left[h_{3}(i)\ \log\nu(t_{i})-\nu(t_{i})\right]{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\frac{|}{|}}$
@@
where $h_{3}(i)$ and $h_{3}^{\text{Po-214}}(i)$ are the observed event counts in phase III in time bin $i$ (with central time $t_{i}$) with the quality cut and 214Po cut applied, respectively. This model is fit to the data using a Markov Chain Monte Carlo (MCMC). Specifically, the python package emcee is used with the “stretch-move” Metropolis-Hastings algorithm [107, 108], propagating the 100 random walkers for $2\times 10^{5}$ steps after an initial burn-in phase of $4\times 10^{4}$ steps. This exceeds the recommendation that the number be at least 50×50$\times$50 × the estimated autocorrelation time [108]. The resulting fits (with statistical uncertainty in the form of random MCMC samples) are shown for each campaign in Fig. 2 in the main text.
@@
Campaign 1 (5.36 $\pm$ 0.62)$\times$10-4 (7.58 $\pm$ 0.66)$\times$10-411footnotemark: 1
(7.07 $\pm$ 5.00)$\times$10-422footnotemark: 2
Campaign 2 (7.55 $\pm$ 0.78)$\times$10-4 (8.63 $\pm$ 0.86)$\times$10-411footnotemark: 1
(4.50 $\pm$ 4.50)$\times$10-422footnotemark: 2
Campaign 3 (8.18 $\pm$ 0.67)$\times$10-4
@@
Figure S3: Top panel Expected decay rates with either the quality cut (light green curves) or 214Po cut (dark green curves) applied, derived from modeling of the phase III data from each campaign (green histograms), as well as the phase IV data (red histograms). Bottom panel Inverse R-values over time, obtained with the quality cut (red curves) and 214Po cut (50%) applied (blue curves), including their respective 90% confidence level (CL) intervals. The vertical dashed line defines the reference time where the rate with quality cut (light green line) is equal to 76.5 Hz.
The rates in phase I with the quality cut are consistent, although campaigns 2 and 3 exhibit a slightly higher rate. This small increase may be attributed to the buildup of 210Po on the inner surface or volume of the SPC due to consecutive injections of radon over time. Similarly, for the 214Po cut, the rates align closely within the uncertainties with a slight increase. The same trend is observed in phase IV when comparing the rates of campaigns 1 and 2 with the quality and 214Po cuts. Nonetheless, the ratio between the rates in phase IV and phase I for campaigns 1 and 2 approaches 1 (see Fig. 3 in the main text), indicating that the radon rate quickly reduces to pre-injection levels in campaigns 1 and 2. This also demonstrates that the performance of the radon trap aligns well with the achievable sensitivity of our experimental setup.
@@
In this section, we describe how the R-value (specifically $R(t)^{-1}$, see Eq. 1 in the main text) is obtained for each measurement campaign. Separate results with and without the 214Po cut are derived by first defining a log likelihood function for the phase IV data of each campaign in each time bin $i$, scaling the phase III decay rate trend by $R(t_{i})^{-1}$ and allowing the background rate $\eta$ to vary as a nuisance parameter:
$\log\mathcal{L}_{R}(i)\ =$ $h_{4}(i)\log\nu(t_{i})-\nu(t_{i}){\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\frac{|}{|}}$
$-\frac{1}{2}\ \left(\frac{\eta-\eta_{\mathrm{obs}}}{\sigma_{\eta}}\right)^{2},$ (S3)
where the latter term is a Gaussian constraint term for the background rate with respect to the observed background rate (from phase I) and its uncertainty ($\eta_{\mathrm{obs}}$ and $\sigma_{\eta}$, respectively). $h_{4}(i)$ is the observed event counts in phase IV in time bin $i$. The expected count rate in the center of each time bin, $\nu(t_{i})$, is equal to the expected rate trend from phase III (see Eq. S1) multiplied by $R(t_{i})^{-1}$, plus $\eta$. For the calculation of $R(t_{i})^{-1}$ with 214Po cuts, we conservatively assume a 50% detection efficiency for 214Po events, due to the possibility that many isotopes may have settled on the detector surface and then emit alphas away from the detector volume.
The best-fit value of $R(t_{i})^{-1}$ is determined in each time bin with the python scipy.optimize.minimize class, with a Nelder-Mead method specifically [92, 93], profiling over the background rate. A $90\%$ confidence level (CL) interval is then calculated using the Feldman-Cousins method [94]. To address significant statistical fluctuations in the data count rate over time, a rolling-average algorithm (over 15 time bins) is applied to the best-fit $R(t)^{-1}$ and 90% CL interval. The above procedure is performed 125 times to include statistical uncertainty on the phase III rate model, varying the model parameters with randomly drawn MCMC samples. This yields an ensemble of slightly different best-fit $R(t_{i})^{-1}$ and corresponding 90% CL intervals for each time bin, of which we take the median result for each.
The $R(t)^{-1}$, along with the corresponding 90% CL intervals, are shown in the bottom panels of Fig. S3 for each campaign. From these results, we determine the final R-value, $R(t)$, at the reference time, which is denoted by vertical dashed lines in Fig. S3. The derived values and their associated 90% CL intervals for either the quality cut or the 214Po cut (50%) are summarized in Table S2.
@@
Campaign 1 9605.64 9376.55
[3465.66, 67957.1] [5286.86, 19292.9]
Campaign 2 10023.5 17057.9
[3234.85, 76679.6] [6449.61, 43815.8]
Campaign 3 5.39002 11.7936
[5.34640, 5.43372] [11.6108, 11.9752]
Table S2: R-values with the 90% CL intervals (lower and upper limits) at the reference time with either quality cut or 214Po cut (50%) applied for each campaign. The 90% lower confidence limit (LCL) is reported as the final result of R-value for each campaign and shown in Fig. 4 in the main text.

### v2

@@ II.1 Quality cut
We apply a quality cut that includes a loose pulse shape discrimination cut to remove non-physical events with characteristically short risetimes [NEWS-G:2017pxg], as well as requiring pulse amplitudes above 1200 ADU (arbitrary digital units). The latter selection removes most of the background events from atmospheric muons, as well as additional electronic noise events. Some degraded alpha events that do not deposit their full kinetic energy also fall below this cut. However, the fraction of affected alpha events remains mostly constant over time, tending to increase slightly as the detector gain decreases. A separate study quantifying the impact of this effect on signal efficiency due to this cut shows a drop of only $\sim$ 1% over time (see Fig. 3.59 in Ref. [Durnford:2024]).
@@
Figure S1: Amplitude distribution of events for approximately 6 hours of data in phase III of campaign 1, recorded by the SPC (black markers). The distribution is modeled with an adaptive-bandwidth Gaussian kernel density estimation [silverman, wang2011bandwidth, kde_article, kde_code] with bootstrapped [efron] statistical uncertainty (red curve and $1\sigma$ shaded band). The inset shows the amplitude peak positions corresponding to the maximum energy deposition of three alpha decays versus the corresponding alpha energies (red markers with error bars), with a linear fit (blue curve and shaded band).
@@
The remaining constant background in phase I for all campaigns (see Fig. 2 in the main text) comes predominantly from the alpha decay of 210Po (5.3 MeV [nndc_206]), originating from radioisotope depositions on the inner surface of the detector material. The SPC, made of stainless steel, does not have particular precautions to prevent 210Po contamination, but can nonetheless serve as a useful low-rate calibration source for this analysis. In phase IV, following the opening of the trap, the rate in campaigns 1 and 2 decreases significantly over several hours, eventually reaching the 210Po alpha background level observed in phase I. For campaign 3, using activated charcoal, the rate remains up to two orders of magnitude higher compared to the rate in phase I and exhibits exponential decay, indicating the continued presence of radon.
@@
In phases II and III, with the injection of radon, the rate increases. By applying the quality cut, we select primarily the full energy alpha decays of  [^2] Rn (5.5 MeV [nndc_218]) and its daughter isotopes, 218Po (6.0 MeV [nndc_214]) and 214Po (7.7 MeV [nndc_210]), which have clear amplitude peaks (above 12,000 ADU) in the recorded data, as shown in Fig. S1 for phase III. The nominal radon trap analysis described in the main text is performed on the data with only this quality cut applied, incorporating events of all three alpha decay isotopes; as these decays are in secular equilibrium, the proportion of events due to radon itself is approximately constant throughout each measurement campaign. Additionally, high energy beta decay events from the  [^2] Rn decay chain (214Bi in particular [nndc_214]) are present in this data selection, but are also in equilibrium with the rest of the decay chain.
@@
To verify the linear energy response of the SPC in this regime, amplitude spectra—such as the one shown in Fig. S1—are characterized with an adaptive-bandwidth Gaussian kernel density estimation [silverman, wang2011bandwidth, kde_article, kde_code], with statistical uncertainties being estimated using a bootstrapping procedure [efron]. The resulting peak amplitudes are well-fit by a linear function of the corresponding alpha decay energies, as shown in the inset of Fig. S1. A peak corresponding to 210Po might be present in this amplitude spectrum, but its rate is more than two orders of magnitude lower than the  [^2] Rn peak with which it overlaps.
@@
In our analysis, we develop a method to select only 214Po events. 214Po decays achieve secular equilibrium with  [^2] Rn due to its short half-life (164 $\upmu$}s [nndc_210]) and the short half-lives of the other intermediate isotopes compared to  [^2] Rn and its subsequent daughter isotopes. By selecting only 214Po events, the signal-to-background ratio is significantly improved, as there is no 210Po background in this amplitude regime.
@@
Due to the large amount of radon remaining in phase IV of campaign 3 (see Fig. 2 in the main text), the 214Po amplitude cut is determined using a similar method as in phases II and III. To obtain a time-dependent 214Po amplitude cut for phase IV, the amplitude minima between the 218Po and 214Po peaks are found in approximately 6-hour data segments. The bootstrap resampling [efron] is then applied to estimate the statistical uncertainty for each point. These points are then fit with a linear function to define the 214Po amplitude cut for phase IV.
@@ II.3 Expected decay rates
To achieve this, a Monte Carlo (MC) simulation of radon decay is produced for a given radon rate in addition to a given background rate, which includes the expected pileup and dead-time effects to reflect the actual conditions of our experiment. Any simulated decays occurring within 1 ms of each other (half the duration of a single event window) are treated as coincident events (resulting in a single recorded event). Simulated decays occurring within 1 to 2 ms of a preceding pulse are considered as lost due to DAQ deadtime. This approach accounts for the DAQ software centering the maximum amplitude of the triggered event trace within its acquisition time window, ensuring that no two pulses can be more than half an event window apart and still be contained in the event window. High-statistics MCs are produced for true radon rates ranging from approximately 30 Hz to 200 Hz, along with the measured 210Po background rate for each campaign described in the main text. Each MC dataset spans a duration longer than the real data from phase III, covering both phase III and phase IV. The resulting log-scale rate trends—separated into single decay event vs. pileup events—are characterized using univariate splines [scipy_spline].
The results of the decay rate MCs are incorporated into a joint model for the observed event rate over time, with and without the application of the 214Po cut. The free parameters of the model define: (i) the true radon decay rate from the discrete sets of values for which MCs are produced; (ii) the overall alpha event signal efficiency with the quality cut applied ($f_{\alpha}$); (iii) signal efficiency of observed 214Po events with the 214Po cut applied ($f_{\text{Po-214}}$); (iv) the fraction of pileup events that pass the quality or 214Po cut ($f_{\text{pileup}}$ and $f_{\text{pileup, Po-214}}$, respectively). Adding the background rates for each data selection ($\eta$ and $\eta_{\text{Po-214}}$), the expected event rates over time $\nu$, with either the quality or 214Po cut, are given as
$\nu$ $(t)=f_{\alpha}\left[10^{y_{\text{Rn-222}}(t)}+10^{y_{\text{Po-218}}(t)}+10^{y_{\text{Po-214}}(t)}\right]{\color[rgb]{1,1,1}\frac{|}{|}}$
$+f_{\text{pileup}}\times 10^{y_{\text{pileup}}(t)}+\eta,{\color[rgb]{1,1,1}\frac{|}{|}}$
$\nu$ $_{\text{Po-214}}(t)=f_{\text{Po-214}}\times 10^{y_{\text{Po-214}}(t)}{\color[rgb]{1,1,1}\frac{|}{|}}$
@@
$\log$ $\mathcal{L}\ =\sum_{i}\left[h_{3}(i)\ \log\nu(t_{i})-\nu(t_{i})\right]{\color[rgb]{1,1,1}\frac{|}{|}}$
@@
where $h_{3}(i)$ and $h_{3}^{\text{Po-214}}(i)$ are the observed event counts in phase III in time bin $i$ (with central time $t_{i}$) with the quality cut and 214Po cut applied, respectively. This model is fit to the data using a Markov Chain Monte Carlo (MCMC). Specifically, the python package emcee is used with the “stretch-move” Metropolis-Hastings algorithm [GW10, emcee], propagating the 100 random walkers for $2\times 10^{5}$ steps after an initial burn-in phase of $4\times 10^{4}$ steps. This exceeds the recommendation that the number be at least $50\times$ the estimated autocorrelation time [emcee]. The resulting fits (with statistical uncertainty in the form of random MCMC samples) are shown for each campaign in Fig. 2 in the main text.
@@
[…]
@@
Figure S3: Top panel Expected decay rates with either the quality cut (light green curves) or 214Po cut (dark green curves) applied, derived from modeling of the phase III data from each campaign (green histograms), as well as the phase IV data (red histograms). Bottom panel The 90% confidence level (CL) intervals of the inverse R-values over time, obtained with the quality cut (red curves) and 214Po cut (50%) applied (blue curves). The vertical dashed line defines the reference time where the extrapolated rate with quality cut (light green line) is equal to 76.5 Hz.
The rates in phase I with the quality cut are nearly consistent, with campaigns 2 and 3 each exhibiting a slightly higher rate than the previous campaign. This small increase may be attributed to the buildup of 210Po on the inner surface or volume of the SPC due to consecutive injections of radon over time. Based on the expected buildup and the lifetime of 210Po, a small increase in rate—on the order of 0.01 Hz—is expected, which aligns well with the observed data. Similarly, for the 214Po cut, the rates align closely within the uncertainties with a slight increase, likely due to the presence of additional 210Po, contributing to more mis-reconstructed events. The same trend is observed in phase IV when comparing the rates of campaigns 1 and 2 with the quality and 214Po cuts. Nonetheless, the ratio between the rates in phase IV and phase I for campaigns 1 and 2 approaches 1 (see Fig. 3 in the main text), indicating that the radon rate quickly reduces to pre-injection levels in campaigns 1 and 2. This also demonstrates that the performance of the radon trap aligns well with the achievable sensitivity of our experimental setup.
@@
In this section, we describe how the R-value (specifically $R(t)^{-1}$, see Eq. 1 in the main text) is obtained for each measurement campaign. As mentioned in the main text, the observed count rate in phase IV is fit with a generic model consisting of a exponential function plus a constant component, in order to smooth–over statistical fluctuations and remove dependence on the binning of the counts over time. This fit was done with an MCMC following the same procedure followed for the expected rate trend described in section S2. Separate results with and without the 214Po cut are derived by first defining a log–likelihood function for the phase IV model of each campaign, at central time of each bin $i$, scaling the phase III decay rate trend by $R(t_{i})^{-1}$ and allowing the background rate $\eta$ to vary as a nuisance parameter:
$\log\mathcal{L}_{R}(i)\ =\$$h_{4}(i)\log\nu(t_{i})-\nu(t_{i})-\frac{1}{2}\ \left(\frac{\eta-\eta_{\mathrm{obs}}}{\sigma_{\eta}}\right)^{2},$ (S3)
where the latter term is a Gaussian constraint term for the background rate with respect to the observed background rate (from phase I) and its uncertainty ($\eta_{\mathrm{obs}}$ and $\sigma_{\eta}$, respectively). $h_{4}(i)$ is the observed count rate model for phase IV in time bin $i$. The expected count rate in the center of each time bin, $\nu(t_{i})$, is equal to the expected rate trend from phase III (see Eq. S1) multiplied by $R(t_{i})^{-1}$, plus $\eta$. For the calculation of $R(t_{i})^{-1}$ with 214Po cuts, we conservatively assume a 50% detection efficiency for 214Po events, due to the possibility that many isotopes may have settled on the detector surface and then emit alphas away from the detector volume.
A $90\%$ confidence level (CL) interval is calculated for each time bin using the Feldman-Cousins method [Feldman:1997qc]. This is repeated for 500 randomly–drawn MCMC samples from the fit of the phase III rate model and the phase IV observed rate model to include statistical uncertainty for both fits. This yields an ensemble of slightly–varying 90% CL intervals for $R(t_{i})^{-1}$ for each time bin, of which we take the median result for each.
The $R(t)^{-1}$ 90% CL intervals are shown in the bottom panels of Fig. S3 for each campaign. From these results, we determine the final R-value, $R(t)$, at the reference time, which is denoted by vertical dashed lines in Fig. S3. The derived values and their associated 90% CL intervals for both cuts are summarized in Table S2.
@@
Campaign 1 [4006.9, inf] [5684.9, 28337]
[…]
Campaign 3 [5.4086, 5.5079] [11.395, 11.886]
Table S2: 90% CL intervals (lower and upper limits) of R-values at the reference time with either quality cut or 214Po cut (50%) applied for each campaign. The 90% lower confidence limit (LCL) is reported as the final result for each campaign and shown in Fig. 4 in the main text.